# TokenLight: Precise Lighting Control in Images using Attribute Tokens

### 16 Apr 2026 v1

@@ 3.3 Modeling Lighting Edits with Diffusion Model
The model is trained using a flow-matching objective [34, 35], which learns a time-dependent velocity field that transports noisy latents toward data latents along straight-line paths. Conditioning is provided by the input image $I$ and its associated lighting edit $\Delta L$. Given a data latent $\mathbf{X}\!\sim\!p(\cdot\mid I,\Delta L)$ and base noise $\boldsymbol{\epsilon}\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{I})$, we form the interpolant $\mathbf{z}_{\tau}=(1-\tau)\,\boldsymbol{\epsilon}+\tau\,\mathbf{X}$, where $\tau\in[0,1]$. The network $u_{\theta}(\mathbf{z}_{\tau},\tau,I,\Delta L)$ is trained to predict the target velocity $\mathbf{X}-\boldsymbol{\epsilon}$, minimizing:
@@ 4.1 Quantitative Evaluation
To measure how well the model distinguishes nearby light positions, we create a confusion matrix $M\in\mathbb{R}^{32\times 32}$, where $M_{ij}$ denotes the error between the prediction conditioned on light position $j$ and the ground truth (GT) rendered at light position $i$. $M$ captures both accuracy (low diagonal error) and sensitivity (off-diagonal error increasing with trajectory distance). In Tab. 3, we report $A$ (lower is better) and $B/A$ (higher is better), where $A,B$ are:
@@
In Fig. 5, we show results along an example trajectory and the confusion maps. We include a GT vs. GT reference. Our method achieves lower $A$, higher $B/A$, and confusion maps closer to GT than Neural Gaffer, indicating more precise spatial light control.
@@ Appendix B Additional Implementation Details
To support intensity, color, spatial, and diffuse controls, we generate supervision on-the-fly during data loading by combining an ambient render $I$, an on-light render $O$, and a tone-mapping operator $\mathbf{T}(\cdot)$.
@@
For spatial lighting, we follow a parallel construction. The ambient image $I$ is given by the environment-map render, while the controllable source is a point light with a sampled 3D position. The render $O$ corresponds to this single point light, and we again form the relit image as
@@
For diffuse-spread supervision, we fix the light position and render multiple spread levels. Let $A$ denote the ambient render, $O_{1}$ and $O_{2}$ the on-light renders with different spread parameters, and $I$ and $I_{r}$ the corresponding tone-mapped input and target images. Training pairs hold intensity, color, and location constant while varying only the spread parameter:
@@
The difference in spread between $O_{1}$ and $O_{2}$, $d_{g}$ conditions the model. Note that in order to ensure that perceived differences in shadow softness only stem from $O_{1}$ and $O_{2}$ we use constant ambient lighting for $A$, avoiding environment map renders as they can introduce strong directional shadows (e.g., from sunlight) that would appear in both $I$ and $I_{r}$, breaking the intended spread-only supervision, since such fixed shadows would confound the effect of changing only the diffuse-spread parameter.
In all three cases, the network predicts $I_{r}$, given $I$ and lighting-edit attributes.
@@ Appendix F Independent control of multiple lights
Extending the single-light formulation in Sec. 3.1, we construct supervision by summing the contributions of individual lights. Let $O_{i}$ denote the render corresponding to the $i$-th light at position $\mathbf{p}_{i}$ with sampled parameters $(\mathbf{c}_{i},\lambda_{i})$, and let $I$ denote the ambient render. The relit image is given by

### 17 Apr 2026 v2

@@ 3.3 Modeling Lighting Edits with Diffusion Model
The model is trained using a flow-matching objective [34, 35], which learns a time-dependent velocity field that transports noisy latents toward data latents along straight-line paths. Conditioning is provided by the input image II and its associated lighting edit $\Delta L$. Given a data latent $\mathbf{X}\!\sim\!p(\cdot\mid I,\Delta L)$ and base noise $\boldsymbol{\epsilon}\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{I})$, we form the interpolant $\mathbf{z}_{\tau}=(1-\tau)\,\boldsymbol{\epsilon}+\tau\,\mathbf{X}$, where $\tau\in[0,1]$. The network $u_{\theta}(\mathbf{z}_{\tau},\tau,I,\Delta L)$ is trained to predict the target velocity $\mathbf{X}-\boldsymbol{\epsilon}$, minimizing:
@@ 4.1 Quantitative Evaluation
To measure how well the model distinguishes nearby light positions, we create a confusion matrix $M\in\mathbb{R}^{32\times 32}$, where $M_{ij}$ denotes the error between the prediction conditioned on light position $j$ and the ground truth (GT) rendered at light position ii. $M$ captures both accuracy (low diagonal error) and sensitivity (off-diagonal error increasing with trajectory distance). In Tab. 3, we report AA (lower is better) and $B/A$ (higher is better), where $A,B$ are:
@@
In Fig. 5, we show results along an example trajectory and the confusion maps. We include a GT vs. GT reference. Our method achieves lower AA, higher $B/A$, and confusion maps closer to GT than Neural Gaffer, indicating more precise spatial light control.
@@ Appendix B Additional Implementation Details
To support intensity, color, spatial, and diffuse controls, we generate supervision on-the-fly during data loading by combining an ambient render II, an on-light render $O$, and a tone-mapping operator $\mathbf{T}(\cdot)$.
@@
For spatial lighting, we follow a parallel construction. The ambient image II is given by the environment-map render, while the controllable source is a point light with a sampled 3D position. The render $O$ corresponds to this single point light, and we again form the relit image as
@@
For diffuse-spread supervision, we fix the light position and render multiple spread levels. Let AA denote the ambient render, $O_{1}$ and $O_{2}$ the on-light renders with different spread parameters, and II and $I_{r}$ the corresponding tone-mapped input and target images. Training pairs hold intensity, color, and location constant while varying only the spread parameter:
@@
The difference in spread between $O_{1}$ and $O_{2}$, $d_{g}$ conditions the model. Note that in order to ensure that perceived differences in shadow softness only stem from $O_{1}$ and $O_{2}$ we use constant ambient lighting for AA, avoiding environment map renders as they can introduce strong directional shadows (e.g., from sunlight) that would appear in both II and $I_{r}$, breaking the intended spread-only supervision, since such fixed shadows would confound the effect of changing only the diffuse-spread parameter.
In all three cases, the network predicts $I_{r}$, given II and lighting-edit attributes.
@@ Appendix F Independent control of multiple lights
Extending the single-light formulation in Sec. 3.1, we construct supervision by summing the contributions of individual lights. Let $O_{i}$ denote the render corresponding to the $i$-th light at position $\mathbf{p}_{i}$ with sampled parameters $(\mathbf{c}_{i},\lambda_{i})$, and let II denote the ambient render. The relit image is given by